OpenStudy (anonymous):

Find the derivative of y = x^2 / 3x-1 with respect to 'x' at x=2

OpenStudy (anonymous):

quotient rule for this one \[(\frac{f}{g})^'=\frac{gf'-fg'}{g^2}\] with \[f(x)=x^2,f'(x)=2x,g(x)=3x-1,g'(x)=3\]